OpenStudy (anonymous):

Solve A = 1/2p + t for p in terms of A and t.

OpenStudy (ajprincess):

1.Subtract t both sides. 2.Multiply by 2 both sides. Can u do nw?

OpenStudy (ajprincess):

@lukelebeau can u do?

OpenStudy (anonymous):

p = A – 2t??

OpenStudy (ajprincess):

hmm no. A=1/2p+t A-t=1/2p 2(A-t)=p 2A-2t=p

OpenStudy (anonymous):

ohh okay i see thanks soo much!!

OpenStudy (ajprincess):

yw.:)